# HAFixAgent: History-Aware Automated Program Repair Agent

### v2

@@ 3.2.2. Bug Category Terminology
To categorize bugs based on the structure of their bug fix commits, we first define a hunk as a single, contiguous block of edits (i.e., additions, deletions, or modifications) applied to a specific region in the source code, similar to earlier work (nashid2025characterizing). A patch file can contain one or more such hunks. Furthermore, following prior work on bug categorization (nashid2025characterizing; bouzenia2024repairagent; jiang2023impact; xia2024automated), we classify all 854 bugs into four mutually exclusive categories based on the number and location of these hunks:
@@
Unlike prior work that either nests SL within SH (jiang2023impact) or collapses SFMH and MFMH into one multi-hunk class (nashid2025characterizing), we keep all four categories separate for finer granularity.
Figure 1. Distribution of the number of unique blame commits across different bug categories on Defects4J.
Table 1. Blame availability by bug category in Defects4J, assuming perfect fault localization. Percentages in the Blameable or Blameless columns are within-category; Total shows the share relative to the full dataset.
Category Blameable Blameless Total
SL 140 83.8 27 16.2 167 19.6
SH 70 52.2 64 47.8 134 15.7
SFMH 294 69.2 131 30.8 425 49.8
MFMH 103 80.5 25 19.5 128 15.0
Total 607 71.1 247 28.9 854 100.0
@@ 3.3. Results
Historical blame commit information is broadly available, even for complex bugs. As shown in Table 1, the majority (71.1%) of bugs in Defects4J are blameable, i.e., have at least one blame commit across all buggy lines. Notably, complex multi-hunk bugs (SFMH and MFMH), which account for 64.8% of the entire dataset, maintain high history availability: 69.2% of SFMH bugs and 80.5% of MFMH bugs are blameable. Meanwhile, SL has 140 of 167 blameable (83.8%) and SH has 70 of 134 (52.2%). SFMH is the largest subset at 49.8% of all bugs.
These numbers indicate that historical blame data is commonly available across categories, and hence can be used to ground the reasoning of an APR technique. That said, for the almost thirty percent for which no blame commit can be determined, a workaround would need to be provided.
All but three of the blameable Defects4J bugs have exactly one (!) unique blame commit. Figure 1 shows the distribution of unique blame commits across all bugs. Contrary to expectations, 70.7% (604 out of 854) of all bugs map to exactly one unique blame commit To validate this finding, we manually inspected a sample of multi-hunk cases and confirmed that their blameable lines did indeed trace back to a single, common commit. While one would expect multi-hunk bugs to require at least one separate blame commit per hunk, it turns out that the (blame) history of buggy hunk lines overlaps almost exactly. Those blame commits could still change other lines in addition to the buggy lines (potentially diluting blame context for APR), but the need to add context for only one blame commit suggests that adapting this heuristic to more complex bug categories could be feasible!
In contrast, 28.9% (247 out of 854) have no blame commits (the add-only, blameless cases), and only 0.4% (3 out of 854) of the bugs have two or more git blame commits. The same pattern holds within each category, for example, SL has 140 bugs with one blame commit and 27 with none, and MFMH has 102 bugs with one blame commit and 25 with none.
Summary for Preliminary Study (RQ0): (1) Blame commits can be identified for 71.1% of bugs (when assuming perfect fault localization). (2) Even for complex bug types, blame history typically is concentrated into one commit, as 70.7% of Defects4J bugs have exactly one unique blame commit.
@@ 4. HAFixAgent
    the agent has access to a git blame tool, inspired by the findings in Section 3, and leverages the single and unique blame commit as the primary source of historical context.
@@
Unlike agent-based APR systems such as RepairAgent (bouzenia2024repairagent), which employ pre-defined Python tool APIs, HAFixAgent is intentionally minimal and history-aware. It treats repository history from previous snapshots as an important source of context that augments the current code snapshot, and it operates within a compact, guarded loop over a small set of standard bash tools (e.g., grep, sed, find). This design choice is critical: it minimizes ambiguity from complex and high-level tool interactions, allowing us to more directly attribute performance gains to the quality of the historical context rather than to sophisticated tool exploration. Additionally, the design is consistent with recent commercial coding assistants such as Claude Code (anthropic_claude_code_docs_2025) and OpenAI Codex (openai_codex_2025).
@@ 4.1.2. History Extractor
The History Extractor module (step 2 in Figure 2) is responsible for finding a relevant blame commit for every bug. We use the buggy lines identified through the fault localization metadata to obtain the blame commit, since the historical context in that blame commit may reveal the root causes of the bug (sliwerski2005changes; hassan2006mining; kamei2012large). The module handles blameable and blameless bugs differently:
  • Blameable bugs. For bugs involving modified or deleted lines, we execute git blame on all identified lines and collect the union of all unique commit hashes. This process is deterministic rather than a random selection of one line. As demonstrated in our preliminary study (Section 3), this union is nearly always a single commit: 70.7% of all bugs (except three blameable bugs) map to exactly one unique blame commit.
@@ 4.1.3. Historical Context
Once the blame commits are determined by the History Extractor, step 3 then mines relevant historical context from the identified blame commits by leveraging one of the three top-performing historical heuristics from our HAFix approach (shi2025hafix) (which focused on SL bugs only):
@@ 4.2. Agent Execution Loop
HAFixAgent’s execution loop follows an observe-act-feedback paradigm, inspired by frameworks like ReAct (yao2023react) and those used in recent software engineering agents (yang2024sweagent). The loop proceeds with the following steps: 5 the LLM receives the current state (including the full conversation history of prior actions and observations); 6 the agent parses the model output to extract a tool invocation (i.e., a bash command, which the system prompt constrains the model always to provide) as the next action; 7 the command runs in a containerized tool sandbox that exposes repair utilities and dataset tools (e.g., search, edit and compile); 8 stdout and stderr are collected from the output of command execution as feedback; 9 compilation and test outcomes are captured and parsed; 10 the loop terminates on success or when a termination guard (e.g., step limit, cost limit, or timeout) is reached; otherwise, $1$ the observation is appended to the conversation history and fed back to the LLM for the next reasoning step.
@@ 5.1. Prompt Context Preparation
To prepare the non-history context as defined in Section 4.1.1, we need bug metadata such as bug reports, failing tests, and fault locations. We mine bug reports from issue tracker links provided by Defects4J. For 18 cases from the Chart project without a clear bug issue link, we use the fixing commit message as the bug description, and to prevent potential data leakage, we manually verified these commit messages to ensure they do not contain explicit fix instructions. Similar to the study (nashid2025characterizing), we collect the failing tests provided by Defects4J. Consistent with common practices in the field of program repair (yang2024sweagent; fan2023automated; bouzenia2024repairagent; xia2022less; ye2024iter; nashid2025characterizing; ehsani2025bug), we assume perfect fault localization and provide fault localization context, which is derived from the developer’s bug fix (following the process in Section 3.2), directly into the agent’s context. Specifically, similar to prior work (bouzenia2024repairagent), this context consists of all file(s) and line(s) that need to be edited to fix the bug.
@@
Motivated by the preliminary study in Section 3, which shows that 71.1% of bugs are blameable and most map to a single unique blame commit, we collect history per bug as follows:
@@
  • Blameable with multiple commits: Only three bugs fall in this case, we use the LLM-as-a-judge (the same LLM as HAFixAgent) to select the single most relevant commit. The model is prompted with the set of blameable lines and asked to choose the line which is most likely related to the bug’s root cause.
@@ 5.2. Baselines
We compare HAFixAgent with two recent APR techniques as baselines, i.e., RepairAgent (bouzenia2024repairagent) and BIRCH-feedback (nashid2025characterizing).
RepairAgent.
RepairAgent is an autonomous LLM-based agent equipped with pre-defined API tooling. The original study evaluated it on 835 bugs (from Defects4J v1.2 and v2), reporting 186 plausible (test-passing) and 164 correct (manual semantic check) fixes. Our comparisons are performed on the 829-bug overlap with our v3 set. For our aggregate effectiveness comparison, we compare our test-passing patches against their reported 186 plausible fixes, as these metrics are equivalent. For our fine-grained, per-category analysis, we are limited to their 164-bug correct set, as bug IDs (which are required for categorization) were only released for this set instead of the plausible bug set. [^2] https://github.com/sola-st/RepairAgent/issues/18
BIRCH-feedback.
BIRCH-feedback targets multi-hunk repairs. We compare results on the 371 overlapping multi-hunk bugs reported in that study and use their best-performing configuration with feedback as the baseline, which we refer to as BIRCH-feedback.
@@ 5.3. Metrics
In line with previous studies (chen2021evaluating; parasaram2024fact; ehsani2025bug; nashid2025characterizing; silva2025repairbench), we report Plausible@1 to evaluate the correctness based on whether the generated patch passes all test cases. Formally:
@@
Complementing Plausible@1, which measures the pass rate as a proportion of all bugs), we also report #Pass (the absolute count of bugs passing the test suite), similar to prior work (nashid2025characterizing). To quantify the unique contribution of diverse historical context, we additionally report #Unique Pass, to show the number of bugs solved by a configuration using historical context but not by its non-history counterpart. We visualize overlaps using Venn diagrams to illustrate multi-way complementarity.
For efficiency, we report the average number of agent steps per bug, the average cost per bug in USD computed from token usage, and the cost stratified by outcome (successful versus failed fixes). To assess the statistical significance of differences in cost and number of steps, we compare the distributions across the four context configurations. Since the data is paired (same bugs, different configurations) and not assumed to be normally distributed, we employ the Friedman test (friedman1937use), a non-parametric test for multiple comparisons. If the Friedman test is significant ($p<0.05$), we conduct pairwise Wilcoxon signed-rank tests (wilcoxon1992individual) for post-hoc analysis to compare each history-aware configuration against the non-history configuration. To control for multiple comparisons (three pairs), we apply the Bonferroni correction (dunn1961multiple) and set the significance threshold $\alpha$.
@@ 5.4. Implementation
HAFixAgent is implemented in Python and builds on mini-swe-agent (yang2024sweagent) for the execution loop, which aligns with our lightweight design. While mini-swe-agent (and swe-agent) is designed naturally to support solving PR issue benchmarks such as SWE-Bench (jimenez2023swebench), it lacks the capability of historical context management (e.g., history localizing and retrieval), the repairing strategies and tools specifically for resolving Defects4J bugs. It is an open-source, lightweight framework with about 100 lines of core loop code, enabling seamless integration of historical context extraction and bash tools orchestration during runtime. We use DeepSeek-V3.2-Exp (liu2024deepseek; deepseek-v3.2-exp-tech-note) via the official API with temperature set to 0.0, based on the official DeepSeek recommendation (deepseek-parameter-settings). Prompts are rendered at runtime with jinja2.
We run HAFixAgent on Ubuntu 20.04, with each bug executing in an isolated Docker container. This sandbox environment is built from the released Defects4J image. For this evaluation, we implement the abstract compile and test tools (described in Section 4.3) as wrappers for the defects4j compile and defects4j test -r commands, respectively. The container is preinstalled with all necessary Defects4J and standard bash tools required by HAFixAgent. For each bug, we cap the agent’s loop at 50 steps, enforce a per-bug cost guard of \$1 USD, and set a 1-hour timeout. After each run, the container is automatically cleaned, while all generated patches and logs are persisted for analysis.
@@ 6.1.1. Motivation
Our prior work (shi2025hafix) shows that injecting historical context improves LLM repair for single-line bugs under simple prompting. History carries heuristic signals that are hard to infer from a static snapshot: the fault introducing change, coevolving functions and files, and often intent hints in commit diffs and messages. RQ0’s finding that most Defects4J bugs, even multi-hunk ones, have only one single historical blame commit, suggests that historical context could also scale to APR techniques for more complex bug types.
This RQ rigorously validates the benefit of blame-derived context on SH, SFMH and MFMH bugs, in the context of our proposed HAFixAgent. To thoroughly validate our approach, we first establish HAFixAgent’s practical viability by benchmarking it against state-of-the-art (SOTA) baselines to determine if our history-aware method is competitive. Second, we isolate the specific impact of our core hypothesis by evaluating the agent across different context configurations (i.e., ”with” versus ”without” history). This analysis is essential to pinpoint and quantify the precise contribution of historical context and to understand how these gains vary by bug category. This motivates the following questions: (a) is HAFixAgent competitive against SOTA baselines? (b) what is the performance impact of adding historical context to an APR agent, and (c) how do these gains vary by bug category?
@@ 6.1.2. Approach
We compare HAFixAgent against two SOTA baselines on shared bug subsets: RepairAgent (829 bugs) and BIRCH-feedback (371 multi-hunk bugs). For RepairAgent, our direct comparison uses the set of bugs they reported as correctly patched, as their full list of plausible bug IDs is unavailable; we also report their published aggregated counts (Section 5.2) for reference. To analyze performance across different bug types, we categorize results by our four bug types (SL, SH, SFMH, MFMH). We pay special attention to the multi-hunk categories (SFMH and MFMH), as their need for coordinated changes may make them more sensitive to historical context. BIRCH-feedback serves as a strong baseline for this specific subset.
In addition, we evaluate the four context configurations defined in Section 4.1 across all 854 Defects4J bugs: HAFixAgent-non-history, HAFixAgent-fn_all, HAFixAgent-fn_pair, and HAFixAgent-fl_diff. Notably, 247 bugs (28.9%) are blameless (fix adds only new lines), precluding standard blame. To enable these blameless bugs to also benefit from historical context, we apply the fallback strategy (defined in Section 4.1.3), ensuring every bug has meaningful historical context.

### v3

@@ 3.2.2. Bug Category Terminology
To categorize bugs based on the structure of their bug fix commits, we first define a hunk as a single, contiguous block of edits (i.e., additions, deletions, or modifications) applied to a specific region in the source code, similar to earlier work (Nashid et al., 2025). A patch file can contain one or more such hunks. Furthermore, following prior work on bug categorization (Nashid et al., 2025; Bouzenia et al., 2025; Jiang et al., 2023; Xia and Zhang, 2024), we classify all 854 bugs into four mutually exclusive categories based on the number and location of these hunks:
@@
Unlike prior work that either nests SL within SH (Jiang et al., 2023) or collapses SFMH and MFMH into one multi-hunk class (Nashid et al., 2025), we keep all four categories separate for finer granularity.
Figure 1. Distribution of unique blame commits per bug category on Defects4J (a) and BugsInPy (b). Defects4J blame is almost entirely concentrated in a single commit, while BugsInPy shows substantial multi-commit blame for complex bug categories.
@@ 3.3. Results
Historical blame commit information is broadly available, even for complex bugs. The majority of bugs in both datasets are blameable, i.e., have at least one blame commit across all buggy lines: 71.1% in Defects4J and 87.4% in BugsInPy. In Defects4J, complex multi-hunk bugs (SFMH and MFMH), which account for 64.8% of the dataset, maintain high history availability: 69.2% of SFMH and 80.5% of MFMH bugs are blameable. SH has the lowest rate at 52.2%, as single-hunk bugs more frequently are pure insertions. In BugsInPy, blame availability is even higher across all categories, with SL at 92.7%, SFMH at 91.6%, and MFMH at 92.3%, while SH remains the lowest at 73.8%, consistent with the Defects4J pattern. These results indicate that historical blame data is commonly available across bug categories and languages, supporting history-aware APR techniques. For bugs with no blame commit (28.9% in Defects4J, 12.6% in BugsInPy), a workaround would be required.
Blame concentration differs sharply across languages: Defects4J is usually single-commit, whereas BugsInPy is often multi-commit. Figure 1(a) shows that 70.7% (604 out of 854) of Defects4J bugs map to exactly one unique blame commit, and only 0.4% (3 out of 854) have two or more. Manual inspection confirmed that multi-hunk bugs typically trace back to a single common commit, as their buggy lines largely overlap in history. In contrast, BugsInPy blame is more scattered, with 51.1% (256 out of 501) of bugs mapping to a single blame commit and 36.3% (182 out of 501) having two or more. Multi-commit proportion rises with bug complexity: 0% for SL, 14.3% for SH, 49.0% for SFMH, and 71.4% for MFMH. Among blameable bugs, only 58.4% in BugsInPy are single-commit, compared to 99.5% in Defects4J. This contrast likely reflects differences in development practices: Python projects in BugsInPy tend to have finer-grained commits touching individual lines across sessions, while Defects4J’s Java projects more often introduce related changes in atomic commits.
@@ 4. HAFixAgent
    the agent has access to a git blame tool, inspired by the findings in Section 3, and leverages the blame commit as the primary source of historical context. When multiple blame commits exist, an LLM-based judge selects the most relevant one.
@@
Unlike agent-based APR systems such as RepairAgent (Bouzenia et al., 2025), which employ pre-defined Python tool APIs, HAFixAgent is intentionally minimal and history-aware. It treats repository history from previous snapshots as an important source of context that augments the current code snapshot, and it operates within a compact, guarded loop over a small set of standard bash tools (e.g., grep, sed, find). This design choice is critical: it minimizes ambiguity from complex and high-level tool interactions, allowing us to more directly attribute performance gains to the quality of the historical context rather than to sophisticated tool exploration. Additionally, the design is consistent with recent commercial coding assistants such as Claude Code (Anthropic, 2025) and OpenAI Codex (OpenAI, 2025b).
@@ 4.1.2. History Extractor
The History Extractor module (step 2 in Figure 2) finds a relevant blame commit for every bug. We use the buggy lines identified through the fault localization metadata to obtain the blame commit, since the historical context in that blame commit may reveal the root causes of the bug (Sliwerski et al., 2005; Hassan, 2006; Kamei et al., 2013). The module handles blameable and blameless bugs differently:
  • Blameable bugs. For bugs involving modified or deleted lines, we execute git blame on all identified lines and collect the union of all unique commit hashes. As demonstrated in our preliminary study (Section 3), this union is often a single commit (99.5% of blameable Defects4J bugs), though BugsInPy bugs more frequently have multiple blame commits (41.6%). When multiple commits exist, we use an LLM-as-a-judge to select the most relevant one (detailed in Section 5.1).
@@ 4.1.3. Historical Context
Once the blame commits are determined by the History Extractor, step 3 then mines relevant historical context from the identified blame commits by leveraging one of the three top-performing historical heuristics from the HAFix approach (Shi et al., 2025) (which focused on SL bugs only):
@@ 4.2. Agent Execution Loop
HAFixAgent’s execution loop follows an observe-act-feedback paradigm, inspired by frameworks like ReAct (Yao et al., 2023) and those used in recent software engineering agents (Yang et al., 2024). The loop proceeds with the following steps: 5 the LLM receives the current state (including the full conversation history of prior actions and observations); 6 the agent parses the model output to extract a tool invocation (i.e., a bash command, which the system prompt constrains the model always to provide) as the next action; 7 the command runs in a containerized tool sandbox that exposes repair utilities and dataset tools (e.g., search, edit and compile); 8 stdout and stderr are collected from the output of command execution as feedback; 9 compilation and test outcomes are captured and parsed; 10 the loop terminates on success or when a termination guard (e.g., step limit, cost limit, or timeout) is reached; otherwise, $1$ the observation is appended to the conversation history and fed back to the LLM for the next reasoning step.
@@ 5.1. Prompt Context Preparation
To prepare the non-history context as defined in Section 4.1.1, we need bug metadata such as bug reports, failing tests, and fault locations. For Defects4J, we mine bug reports from issue tracker links provided by the benchmark. For 18 cases from the Chart project without a clear bug issue link, we use the fixing commit message as the bug description, and to prevent potential data leakage, we manually verified these commit messages to ensure they do not contain explicit fix instructions. For BugsInPy, we intentionally omit bug descriptions to avoid data leakage from commit messages, relying on fault locations and failing tests only. Failing tests are collected from each benchmark’s test framework. Consistent with common practices in the field of program repair (Yang et al., 2024; Fan et al., 2023; Bouzenia et al., 2025; Xia and Zhang, 2022; Ye and Monperrus, 2024; Nashid et al., 2025; Ehsani et al., 2025), we assume perfect fault localization and provide fault localization context, which is derived from the developer’s bug fix (following the process in Section 3.2), directly into the agent’s context.
@@
Motivated by the preliminary study in Section 3, which shows that blame commits are broadly available in both datasets, we collect history per bug as follows:
@@
  • Blameable with multiple commits: In Defects4J, only three bugs fall in this case; in BugsInPy, 182 bugs have two or more blame commits. We use the LLM-as-a-judge (the same LLM as HAFixAgent) to select the single most relevant commit. The model is prompted with the set of blameable lines and asked to choose the line which is most likely related to the bug’s root cause.
@@ 5.2. Baselines
We compare HAFixAgent with two recent APR techniques as baselines, i.e., RepairAgent (Bouzenia et al., 2025) and BIRCH-feedback (Nashid et al., 2025). RepairAgent is an autonomous LLM-based agent equipped with pre-defined API tooling. The original study used GPT-3.5-turbo and evaluated on 835 bugs (from Defects4J v1.2 and v2), reporting 186 plausible and 164 correct fixes. We replicate RepairAgent on DeepSeek-V3.2-Exp and evaluate on all 854 Defects4J bugs, using the plausible metric (test-passing patches) for a direct, same-metric comparison. BIRCH-feedback targets multi-hunk repairs. The original study used o4-mini and reported results on 371 multi-hunk bugs. We replicate their best-performing configuration with feedback on DeepSeek-V3.2-Exp and evaluate on the same 371 overlapping multi-hunk bugs (244 SFMH + 127 MFMH).
To eliminate the model confound and ensure a fair comparison, we replicate both baselines on the same LLM used by HAFixAgent (DeepSeek-V3.2-Exp), rather than relying on their originally reported numbers obtained with different models.
@@ 5.3. Metrics
In line with previous studies (Chen et al., 2021a; Parasaram et al., 2025; Ehsani et al., 2025; Nashid et al., 2025; Silva and Monperrus, 2025), we report Plausible@1 to evaluate the correctness based on whether the generated patch passes all test cases. Formally:
@@
Complementing Plausible@1, which measures the pass rate as a proportion of all bugs, we also report #Pass (the absolute count of bugs passing the test suite), similar to prior work (Nashid et al., 2025). To quantify the unique contribution of diverse historical context, we additionally report #Unique Pass, to show the number of bugs solved by a configuration using historical context but not by its non-history counterpart. We visualize overlaps using Venn diagrams to illustrate multi-way complementarity.
For efficiency, we report the average number of agent steps per bug, the average cost per bug in USD computed from token usage, and the cost stratified by outcome (successful versus failed fixes). To assess the statistical significance of differences in cost and number of steps, we compare the distributions across the four context configurations. Since the data is paired (same bugs, different configurations) and not assumed to be normally distributed, we employ the Friedman test (Friedman, 1937), a non-parametric test for multiple comparisons. If the Friedman test is significant ($p<0.05$), we conduct pairwise Wilcoxon signed-rank tests (Wilcoxon, 1992) for post-hoc analysis to compare each history-aware configuration against the non-history configuration. To control for multiple comparisons (three pairs), we apply the Bonferroni correction (Dunn, 1961) and set the significance threshold $\alpha$.
@@ 5.4. Implementation
HAFixAgent is implemented in Python and builds on mini-swe-agent (Yang et al., 2024) for the execution loop, which aligns with our lightweight design. While mini-swe-agent (and swe-agent) is designed naturally to support solving PR issue benchmarks such as SWE-Bench (Jimenez et al., 2024), it lacks the capability of historical context management (e.g., history localizing and retrieval), the repairing strategies and tools specifically for resolving Defects4J bugs. It is an open-source, lightweight framework with about 100 lines of core loop code, enabling seamless integration of historical context extraction and bash tools orchestration during runtime. All experiments (HAFixAgent, RepairAgent, and BIRCH-feedback) use DeepSeek-V3.2-Exp (DeepSeek-AI et al., 2024; DeepSeek-AI, 2025) via OpenRouter with temperature 0.0. Prompts are rendered at runtime with jinja2.
We run HAFixAgent on Ubuntu 20.04, with each bug executing in an isolated Docker container. For Defects4J, the sandbox is built from the released Defects4J image, with the compile and test tools (Section 4.3) wrapping defects4j compile and defects4j test -r. For BugsInPy, we build pre-warmed Docker images per project with the correct Python environment, and the test tool wraps bugsinpy-test. For each bug, we cap the agent’s loop at 50 steps, enforce a per-bug cost guard of \$1 USD, and set a 1-hour timeout. After each run, the container is automatically cleaned, while all generated patches and logs are persisted for analysis.
@@ 6.1.1. Motivation
Prior work (Shi et al., 2025) shows that injecting historical context improves LLM repair for single-line bugs under simple prompting. History carries heuristic signals that are hard to infer from a static snapshot: the fault introducing change, coevolving functions and files, and often intent hints in commit diffs and messages. RQ0’s finding that blame commits are broadly available suggests that historical context could scale to more complex bug types. However, RQ0 also reveals that blame concentration differs across benchmarks: while Defects4J blame is highly concentrated (99.5% single-commit), BugsInPy shows more scattered blame (41.6% multi-commit), raising the question of whether a single-commit strategy generalizes across languages.
This RQ rigorously validates the benefit of blame-derived context on SH, SFMH and MFMH bugs, in the context of our proposed HAFixAgent, across both Java and Python benchmarks. To thoroughly validate our approach, we first establish HAFixAgent’s practical viability by benchmarking it against state-of-the-art (SOTA) baselines on the same LLM to determine if our history-aware method is competitive. Second, we isolate the specific impact of our core hypothesis by evaluating the agent across different context configurations (i.e., with versus without history) on both Defects4J and BugsInPy. This analysis is essential to quantify the precise contribution of historical context, understand how gains vary by bug category, and determine whether the effect generalizes across languages. This motivates the following questions: (a) is HAFixAgent competitive against SOTA baselines on the same LLM? (b) what is the performance impact of adding historical context to an APR agent, and (c) how do these gains vary by bug category and across benchmarks?
@@ 6.1.2. Approach
To ensure a fair comparison, we replicate the two SOTA baselines on the same LLM used by HAFixAgent (DeepSeek-V3.2-Exp), eliminating the model confound present in prior work. We compare HAFixAgent against RepairAgent (Bouzenia et al., 2025) on all 854 Defects4J bugs and BIRCH-feedback (Nashid et al., 2025) on 371 multi-hunk bugs (244 SFMH + 127 MFMH), matching their original scope. To analyze performance across different bug types, we categorize results by our four bug types (SL, SH, SFMH, MFMH). We pay special attention to the multi-hunk categories (SFMH and MFMH), as their need for coordinated changes may make them more sensitive to historical context.
In addition, we evaluate the four context configurations defined in Section 4.1 across all 854 Defects4J bugs and all 501 BugsInPy bugs. The four configurations are HAFixAgent-non-history, HAFixAgent-fn_all, HAFixAgent-fn_pair, and HAFixAgent-fl_diff. Blameless bugs (28.9% in Defects4J, 12.6% in BugsInPy) use the fallback strategy (Section 4.1.3), ensuring every bug has meaningful historical context.